OpenStudy (anonymous):

Can someone help me turn (-5,0) (0,2) in a function equation?

OpenStudy (anonymous):

@AJ01

OpenStudy (anonymous):

find the slope m= (y2-y1)/(x2-x1) then y-y1=m(x-x1)

OpenStudy (anonymous):

Yea I tried I couldnt figure it out...

OpenStudy (anonymous):

m=(2-0)/(0-(-5))=2/5

OpenStudy (anonymous):

y-0=2/5(x-(-5)) y=2/5x+2

OpenStudy (anonymous):

This is what I am doing

OpenStudy (anonymous):

I'm pretty sure I did b right, but I am stuck on c

OpenStudy (anonymous):

do as above

OpenStudy (anonymous):

I graphed (-2,0) (0,2) for a. Then I graphed (-5,0) (0,2) for b. Now I am stuck in c & I am not sure if I did b right.

OpenStudy (anonymous):

find the 2 points from the plot

OpenStudy (anonymous):

What do you mean? I graphed the 2 points in paint...

OpenStudy (anonymous):

So it would be g(x)=2/5x+2?

OpenStudy (anonymous):

graph b

OpenStudy (anonymous):

I did graph b.. I got (-5,0) (0,2)

OpenStudy (anonymous):

the points (0,2) and (-5,0)

OpenStudy (anonymous):

Did you see the screen shot?

OpenStudy (anonymous):

this is the same as above so i am looking to your answer for part a and b...... a is wrong f(x) cannot be negative

OpenStudy (anonymous):



OpenStudy (anonymous):

Okay so C would be g(x)=2/5x+2?

OpenStudy (anonymous):

yes but this if you did part b right!

OpenStudy (anonymous):

Can you check?

OpenStudy (anonymous):



OpenStudy (anonymous):

the plot just shifted by 3 unit x->x-3

OpenStudy (anonymous):

now find 2 points

OpenStudy (anonymous):

Would it be -5 or 1?

OpenStudy (anonymous):

because -2 - 3 is -5

OpenStudy (anonymous):

this one has nothing to do with your first question f(x)=|x+2| x->x-3 y->y so g(x)=|x-3+2| g(x)=|x-1|